200cm^3 of 0.2mol/dm^3 lead trioxoxnitrate(5) are mixed with 200cm^3 of 0.2mol/dm^3 sodium chloride. Calculate the volume of excess reagent. The mass of lead(2) chloride formed.


1
Expert's answer
2021-03-17T06:44:52-0400

2NaCl + Pb(NO3)2 → 2NaNO3 + PbCl2

Mole of Pb(NO3)2 = 0.2 L × 0.2 M = 0.04mole

Mole of NaCl = 0.04 mole

Limiting reagent = NaCl

Excess reagent = Pb(NO3)2


2NaCl = PbCl2

1 mole NaCl = 1/2 mole of PbCl2

1 mole NaCl = 1/2 × 278 g PbCl2

0.04 mole NaCl = 278 × 0.04 /2 g PbCl2

= 5.56g